Express the height of a cylindrical can as a function of the diameter, if 18pi sq.in. of sheet metal is used. H=?


1
Expert's answer
2021-10-29T02:53:42-0400

The total surface area of a solid is the sum of the areas of all of its faces. A cylinder has 3 faces:

the top and bottom circular bases, and the lateral surface of the cylinderю

If the diameter of a circular base is "D," then the area of each base must be:


"A_{base}=\dfrac{\pi D^2}{4}"

"A_{lateral}=2\pi(\dfrac{D}{2})H=\pi DH"

"A_{total}=2A_{base}+A_{lateral}=\dfrac{\pi D^2}{4}+\pi DH"

If "18\pi" sq.in. of sheet metal is used


"\dfrac{\pi D^2}{4}+\pi DH=18\pi"

"H=H(D)=\dfrac{18}{D}-\dfrac{D}{4}"

"H(D)=\dfrac{18}{D}-\dfrac{D}{4} , inches"